Domain and range transformation: Using math induction

T(1)= 1

T(n) = T(n/2) + bnlg n, n > 1
Answer #1

T(1)= 1

T(n) = T(n/2) + bnlg n, n > 1

prove

TI: 2 Tln) = (n/2) + bhlogn, hol P/2) = T (2) = T (7) + 621092 na thus Assume (m) is true thad for pind n=l, niven. is terwe

hope you understand

Happy learning rate well:)